# From Intention To Implementation: Automating Biomedical Research via LLMs

### v2

@@ Abstract
Conventional biomedical research is increasingly labor-intensive due to the exponential growth of scientific literature and datasets. Artificial intelligence (AI), particularly Large Language Models (LLMs), has the potential to revolutionize this process by automating various steps. Still, significant challenges remain, including the need for multidisciplinary expertise, logicality of experimental design, and performance measurements. This paper introduces BioResearcher, the first end-to-end automated system designed to streamline the entire biomedical research process involving dry lab experiments. BioResearcher employs a modular multi-agent architecture, integrating specialized agents for search, literature processing, experimental design, and programming. By decomposing complex tasks into logically related sub-tasks and utilizing a hierarchical learning approach, BioResearcher effectively addresses the challenges of multidisciplinary requirements and logical complexity. Furthermore, BioResearcher incorporates an LLM-based reviewer for in-process quality control and introduces novel evaluation metrics to assess the quality and automation of experimental protocols. BioResearcher successfully achieves an average execution success rate of 63.07% across eight previously unmet research objectives. The generated protocols averagely outperform typical agent systems by 22.0% on five quality metrics. The system demonstrates significant potential to reduce researchers’ workloads and accelerate biomedical discoveries, paving the way for future innovations in automated research systems.
@@ 1 Introduction
Biomedical research is a fundamental driving force behind human development. By uncovering the underlying mechanisms of diseases (Council et al., 2005), biomedical research improves global health, extends life expectancy, and enhances the quality of life. It also fuels economic growth, scientific advancements, and overall societal well-being.
Traditional biomedical research relies heavily on labor-intensive processes like manual data collection, comprehensive literature reviews, complex experimental designs, and extensive data analysis. Although the conventional approach has facilitated notable breakthroughs in disease prevention (Huang et al., 2023; Wang et al., 2023d; Agarwal et al., 2024), diagnosis (Therriault et al., 2024; Zheng et al., 2024; Sepahi et al., 2024), and treatment (Chang et al., 2024; Apperloo et al., 2024; Douvaras et al., 2024), it struggles to keep pace with the data explosion. For example, PubMed now hosts over 37 million citations  [^1] https://pubmed.ncbi.nlm.nih.gov/?Db=pubmed, overwhelming researchers and making it challenging to stay updated with the latest findings. Moreover, traditional research often demands repetitive tasks or interdisciplinary skills, such as coding, hindering research efficiency.
Artificial intelligence (AI) is emerging as a valuable tool in biomedical research, enhancing specific steps in the research pipeline. For instance, novel machine learning models are designed to analyze data and make decisions, including predicting drug-target from compound-protein interactions (Gupta et al., 2021; Zhu et al., 2021; Mak et al., 2023), detecting the presence of cancer from medical images(Bizzo et al., 2021; of Radiology, ESR), estimating patient outcomes from medical history (Li et al., 2023a; Nordin et al., 2023; Malibari, 2023), and so on. The advancement of Large Language Models (LLMs) further supports AI’s role in academic writing (Bom, 2023; Lingard, 2023) and literature summarizing (Sami et al., 2024). However, biomedical research is still time-consuming. The potential for automating the entire biomedical research process remains largely unexplored (Gao et al., 2024), which allows researchers to concentrate on innovation and strategic decision-making.
@@
Secondly, biomedical research is logically complex. On the one hand, it requires a coherent understanding of literature with intricate logical structures, yet existing LLMs perform poorly on critical literature analysis. For example, irrelevant information in lengthy research papers causes catastrophic forgetting of important facts. On the other hand, biomedical research involves breaking down complex problems into logically related sub-tasks. For example, studying pyroptosis in dedifferentiated liposarcoma (DDL) involves multiple interconnected tasks like analyzing the expression variation and genetic changes of PRGs, performing immune infiltration analyses, identifying PRG-related clusters, characterizing the tumor microenvironment within these clusters, and developing a prognostic gene model based on these clusters (Chen et al., 2024). Each task is logically interdependent, making planning and execution by LLMs challenging.
@@
Lastly, it is crucial to measure the performance of the research assistant. From the quality control perspective, assessing the results of intermediate steps ensures that the research assistant provides reliable final output. From the evaluation perspective, detecting the errors of end-to-end responses helps to identify the weaknesses and strengths of different systems and sheds insight for future improvements. Due to the complexity of biomedical research tasks, manually conducting fine-grained evaluation is infeasible; for example, researchers are not patient in labeling errors in an experimental protocol line by line. Current automatic evaluations such as ROUGE (Lin, 2004) and BLEU (Papineni et al., 2002) compute the overlap between the LLM’s generation and the ground-truth answer, which is unavailable. Furthermore, they focus on textual quality while ignoring aspects such as technical completeness and correctness, which are more important in assessing the quality of experiment protocols.
@@
This paper introduces BioResearcher, an intelligent research assistant designed to automate the entire biomedical research process. BioResearcher can take any research objective and conditions the user provides, survey relevant literature, design an appropriate experiment protocol [^2] Our effort primarily focuses on dry lab experiments, which rely on computational methods to conduct bioinformatics analysis. Future research endeavors may extend to wet lab experiments., write programs to implement the protocol and derive meaningful conclusions. BioResearcher develops novel techniques to address the three challenges above.
@@
BioResearcher employs a modular multi-agent architecture based on LLMs to integrate multidisciplinary skills. It comprises four modules: Search, Literature Processing, Experimental Design, and Programming, each containing multiple specialized agents. These agents specialize in distinct tasks, including literature and dataset search, filtering, reports generation from literature, reports analysis, experimental protocol design, dry lab experiment extraction, code writing and execution, and review. The specialization adapts LLMs to different task requirements, e.g., a search agent is more professional in retrieving relevant literature than a general purposed LLM. The collaboration among multiple agents and modules increases the overall performance. As illustrated in Figure 1(b), the Experimental Design module generates a comprehensive experimental protocol detailing specific datasets, methodologies, and standards. This achievement is facilitated by the effective collection, processing, and analysis of relevant literature and datasets by the Search and Literature Processing modules.
@@
BioResearcher adopts a hierarchical learning approach to decompose the complex logical structure. The Literature Processing module standardizes relevant papers into experimental reports to minimize unimportant information and provides analyses. The Experimental Design module then uses the Retrieval-Augmented Generation (RAG) (Gao et al., 2023) technique aided by the analyses to learn knowledge at different levels of granularity, including relevant headings, outlines, and experimental details, thereby facilitating the design of new protocols in a stepwise manner.
@@
  • We introduce the first end-to-end automated system tailored explicitly for biomedical research, significantly reducing researchers’ manual workloads and enhancing efficiency.
@@
  • By exploring the potential of BioResearcher in automating biomedical research, this study lays the groundwork for future innovations in automated research systems, thereby accelerating biomedical discoveries.
@@ 2 Related Work
Artificial intelligence (AI) has improved biomedical applications about processing different textual data (Jin et al., 2024; Li et al., 2023b; Sudarshan et al., 2024). Conventionally, small-scale language models like BERT (Devlin, 2018) were used, but scaling laws indicate that increasing model parameters brings enhanced performance, leading to superior reasoning capability and higher answer accuracy. Therefore, we have seen large amounts of applications based on large language models (LLMs) (Singhal et al., 2023; Zhang et al., 2023; Chen et al., 2023; Bolton et al., 2024; Li et al., 2023c).
2.1 Biomedical LLMs
@@
Large language models (LLMs) are generally pre-trained on vast open-domain corpora but lack domain-specific knowledge. To enhance domain-specific performance, several techniques are used: (1) Fine-tuning optimizes the total or a part of parameters to improve the model performance on a small, specific dataset (Hu et al., 2021; Singhal et al., 2023; Chen et al., 2023; Bolton et al., 2024). (2) Reinforcement learning with human feedback (RLHF) or AI feedback (RLAIF) updates the model parameters to align LLM’s responses with responses from humans or a teacher-model via a reinforcement learning framework (Ouyang et al., 2022; Zhang et al., 2023; Li et al., 2023c). (3) Prompt engineering(White et al., 2023) involves giving instructions or examples to LLM to enforce rules or enhance reasoning(Maharjan et al., 2024; Nachane et al., 2024; Wang et al., 2023b). These methods mainly focus on question-answering tasks, like answering medical consultations (Zhang et al., 2023), taking medical examinations (Singhal et al., 2023) and summarizing clinical reports (Sudarshan et al., 2024). However, LLMs still struggle with complex tasks requiring deeper understanding and reasoning(Gao et al., 2024). For scientific research tasks like generating executable experimental protocols, model responses are much longer and professional and involve intricate, long logical chains. Fine-tuning and RLHF are infeasible due to the lack of instruction data, and simple prompt engineering cannot empower LLM with deep logical reasoning for professional topics.
2.2 LLM-based Agents for Research
@@
LLM-based agents offer significant advantages over the direct use of LLMs, such as actively acquiring information, interacting with environments, and stronger reasoning and planning(Gao et al., 2024). LLMs can function as a single agent by being assigned with specific roles through in-domain fine-tuning (Weng et al., 2024) or role-specific prompts(Boiko et al., 2023; Nori et al., 2023; Park et al., 2023; Wang et al., 2023a; Sudarshan et al., 2024; Fernando et al., 2023; Yang et al., 2023). In contrast, a Multi-Agent System (MAS) comprises multiple LLM-based agents, enabling task completion through various cooperative methods. One approach involves assigning distinct roles to agents, who then reach consensus through negotiation, like discussing clinical diagnosis via multi medical experts(Tang et al., 2024; Fan et al., 2024; Li et al., 2024). However, the consensus can be unreliable due to potential instability and hallucinations from LLMs. Alternatively, MAS can distribute complex tasks into sub-tasks among agents, such as dividing a scientific discovery process into creating idea, experimentation, and writing (Lu et al., 2024). MAS is promising for tasks with complex logical chains and highly detailed requirements. After decomposing tasks and assigning them to agents, we can synthesize the results effectively. Furthermore, employing a professional and rigorous workflow framework helps constrain the agents’ generative processes, ensuring highly reliable outcomes.
@@
Due to the robust performance of LLM-based multi-agent systems (MAS), AI for Research (AI4Research), with MAS at its core, has gained significant attention, particularly in automating scientific workflows. A survey categorizes AI agents into four automation levels based on their proficiency in experimental design and execution (Gao et al., 2024). Level 0 agents can only perform specific predefined tasks (Lim et al., 2023). Level 1 agents can design simple experimental protocols and use in-silico or lab tools (Boiko et al., 2023; Zhou et al., 2023). Level 2 agents, which include our work, are capable of developing rigorous experimental protocols and employing ex-silico tools and statistical methods for hypothesis evaluation (Lu et al., 2024; M. Bran et al., 2024). Level 3 agents, which remain undeveloped, are envisioned to discover new methods and employ diverse techniques to measure biological phenomena.
@@
Current AI4Research systems are infeasible for the biomedical field. Firstly, most systems are designed for computer science (CS) (Weng et al., 2024; Baek et al., 2024; Lu et al., 2024; Su et al., 2024; Yang et al., 2024), where public datasets like OpenReview are available, facilitating training and evaluation. Such extensive review data is lacking in the biomedical domain. Moreover, existing studies emphasize novelty based on literature summaries rather than reliability. Our goal, however, is to design experimental protocols with a focus on reliability and feasibility, necessitating fine-grained literature analysis. Additionally, the focus on experimental design differs across disciplines, further complicating the application of current systems to the biomedical field. For example, in AI, a recent study introduced an AI Scientist (Lu et al., 2024) to enhance and improve existing solutions based on a given task and initial experimental code. Conversely, BioResearcher often designs a series of customized experiments for a new research subject rather than improving existing solutions.
Recently, some studies have proposed AI assistants for specific subfields of biomedical research. For instance, Genesis (Tiukova et al., 2024) focuses on systems biology but requires manual maintenance of a specialized database and the definition of specialized templates, thereby limiting its level of automation. Additionally, CRISPR-GPT (Huang et al., 2024) automates gene editing experiment design with a highly customized workflow, addressing only a segment of the biomedical research process. These systems are challenging to extend to other biomedical research fields. In contrast, our system automates the entire process of biomedical research regardless of the specific field.
@@ 3 BioResearcher
Figure 2: The flowchart of BioResearcher. On the left, the system is divided into four main modules: Search, Literature Processing, Experimental Design, and Programming. The right side illustrates the system’s workflow: it initiates with the Search module that retrieves pertinent literature and datasets based on the user input. The Literature Processing module then converts obtained literature into standardized experimental reports, analyzing them against the research objective, conditions, and requirements. It also works with the Search module to evaluate the applicability of the datasets mentioned in the reports. These processed reports, analyses, and suitable datasets are then sent to the Experimental Design module, which designs an experimental protocol aligned with the research objective. Finally, the Programming module extracts the dry lab experimental tasks from the protocol and generates executable code to perform these tasks. Components underlined are agents based on LLMs.
@@ 3.1 Framework Overview
To emulate the workflow of human researchers, BioResearcher consists of four primary modules: Search, Literature Processing, Experimental Design, and Programming. As shown in Figure 2, the research process begins with the Search module, which comprehends the user input, generates appropriate queries and searches for relevant research papers and datasets from online repositories. The retrieved literature is then filtered, downloaded, and forwarded to the Literature Processing module. Here, each research paper is standardized into an experimental report, and each report is analyzed in light of the user-specified objective, conditions, and requirements. This module also interacts with the Search module to identify the usability of the datasets mentioned in the reports. The processed reports, analyses, and datasets are then forwarded to the Experimental Design module, which constructs an experimental protocol. Finally, the Programming module extracts a sequence of dry lab experiment tasks from the protocol and generates accurate and executable codes for these tasks.

### v4

@@ Abstract
Conventional biomedical research is increasingly labor-intensive due to the exponential growth of scientific literature and datasets. Artificial intelligence (AI), particularly Large Language Models (LLMs), has the potential to revolutionize this process by automating various steps. Still, significant challenges remain, including the need for multidisciplinary expertise, logicality of experimental design, and performance measurements. This paper introduces BioResearcher, the first end-to-end automated system designed to streamline the entire biomedical research process involving dry lab experiments. BioResearcher employs a modular multi-agent architecture, integrating specialized agents for search, literature processing, experimental design, and programming. By decomposing complex tasks into logically related sub-tasks and utilizing a hierarchical learning approach, BioResearcher effectively addresses the challenges of multidisciplinary requirements and logical complexity. Furthermore, BioResearcher incorporates an LLM-based reviewer for in-process quality control and introduces novel evaluation metrics to assess the quality and automation of experimental protocols. BioResearcher successfully achieves an average execution success rate of 63.07% across eight previously unmet research objectives. The generated protocols, on average, outperform typical agent systems by 22.0% on five quality metrics. The system demonstrates significant potential to reduce researchers’ workloads and accelerate biomedical discoveries, paving the way for future innovations in automated research systems.
keywords:
Biomedical Research, AI for Research, Large Language Models, Multi-Agent Systems, Automation
@@ 1 Introduction
Biomedical research is a fundamental driving force behind human development. By uncovering the underlying mechanisms of diseases [1], biomedical research improves global health, extends life expectancy, and enhances life quality. It also fuels economic growth, scientific advancements, and societal well-being.
Traditional biomedical research relies heavily on labor-intensive processes like manual data collection, comprehensive literature reviews, complex experimental designs, and extensive data analysis. Although the conventional approach has facilitated notable breakthroughs in disease prevention [2, 3, 4], diagnosis [5, 6, 7], and treatment [8, 9, 10], it struggles to keep pace with the data explosion. For example, PubMed now hosts over 37 million citations  [^1] https://pubmed.ncbi.nlm.nih.gov/?Db=pubmed, overwhelming researchers and making it challenging to stay updated with the latest findings. Moreover, traditional research often demands repetitive tasks or interdisciplinary skills, such as coding, hindering research efficiency.
Artificial intelligence (AI) is emerging as a valuable tool in biomedical research, enhancing specific steps in the research pipeline. For instance, novel machine learning models are designed to analyze data and make decisions, including predicting drug-target from compound-protein interactions [11, 12, 13], detecting the presence of cancer from medical images [14, 15], estimating patient outcomes from medical history [16, 17, 18], and so on. The advancement of Large Language Models (LLMs) further supports AI’s role in academic writing [19, 20] and literature summarizing [21]. However, biomedical research is still time-consuming. The potential for automating the entire biomedical research process remains largely unexplored [22], which allows researchers to concentrate on innovation and strategic decision-making.
@@
Secondly, biomedical research is logically complex. On one hand, it requires a coherent understanding of literature with intricate logical structures, yet existing LLMs perform poorly on critical literature analysis. For example, irrelevant information in lengthy research papers causes catastrophic forgetting of important facts. On the other hand, biomedical research involves breaking down complex problems into logically related subtasks. For example, studying pyroptosis in dedifferentiated liposarcoma (DDL) involves multiple interconnected tasks like analyzing the expression variation and genetic changes of PRGs, performing immune infiltration analyses, identifying PRG-related clusters, characterizing the tumor microenvironment within these clusters, and developing a prognostic gene model based on these clusters [23]. Each task is logically interdependent, making planning and execution by LLMs challenging.
@@
Lastly, it is crucial to measure the performance of the research assistant. From the quality control perspective, assessing the results of intermediate steps ensures that the research assistant provides reliable final outputs. From the evaluation perspective, detecting the errors of end-to-end responses helps to identify the weaknesses and strengths of different systems and sheds insight for future improvements. Due to the complexity of biomedical research tasks, manually conducting fine-grained evaluation is infeasible; for example, researchers are not patient with labeling errors in an experimental protocol line by line. Current automatic evaluations, such as ROUGE [24] and BLEU [25], compute the overlap between the LLM’s generation and the ground-truth answer, which is unavailable. Furthermore, they focus on textual quality while ignoring aspects such as technical completeness and correctness, which are more important in assessing the quality of experiment protocols.
@@
This paper introduces BioResearcher [^2] Our code and prompts are available at https://github.com/XMUDM/BioResearcher., an intelligent research assistant designed to automate the entire biomedical research process. BioResearcher can take any research objective and conditions the user provides, survey relevant literature, design an appropriate experiment protocol [^3] Our effort primarily focuses on dry lab experiments, which rely on computational methods to conduct bioinformatics analysis. Future research endeavors may extend to wet lab experiments., write programs to implement the protocol and derive meaningful conclusions. BioResearcher develops novel techniques to address the three challenges above.
@@
BioResearcher employs a modular multi-agent architecture to integrate multidisciplinary skills. It comprises four modules: Search, Literature Processing, Experimental Design, and Programming, each containing multiple specialized agents. These agents specialize in distinct tasks, including literature and dataset search, filtering, reports generation from literature, reports analysis, experimental protocol design, dry lab experiment extraction, code writing and execution, and review. The specialization adapts LLMs to different task requirements, e.g., a search agent is more professional in retrieving relevant literature than a general-purpose LLM. The collaboration among multiple agents and modules increases the overall performance. As illustrated in Figure 1(b), the Experimental Design module generates a comprehensive experimental protocol detailing specific datasets, methodologies, and standards. This achievement is facilitated by the effective collection, processing, and analysis of relevant literature and datasets by the Search and Literature Processing modules. Compared with agent systems employing planning agents powered by LLMs to determine agent participation and intervention strategies throughout task execution [26, 27], our system adopts a professionally designed and rigorously structured workflow framework that constrains agent generation processes through systematic procedural constraints, thereby ensuring enhanced stability and reproducibility. Furthermore, our framework diverges from existing human-curated multi-agent systems [28, 29, 30]. In computer science and other fields, current studies often prioritize novelty. However, due to the nature of our biomedical applications, we introduce the Literature Processing module to ensure the reliability and feasibility of our approach in the biomedical domain.
@@
BioResearcher adopts a hierarchical learning approach to decompose the complex logical structure. The Literature Processing module standardizes relevant papers into experimental reports to minimize unimportant information and provides analyses. The Experimental Design module then uses the Retrieval-Augmented Generation (RAG) [26] technique, aided by the analyses, to learn knowledge at different levels of granularity, including relevant headings, outlines, and experimental details, thereby facilitating the design of new protocols in a stepwise manner.
@@
  • We present the first end-to-end automated system designed specifically for biomedical research. At its core is a multi-agent framework powered by LLMs, which decomposes complex research tasks into specialized subtasks. By enabling collaborative execution among domain-specific agents, the system enhances overall performance, significantly reduces manual effort, and improves efficiency.
@@
  • This study explores BioResearcher’s potential to automate biomedical research, paving the way for future innovations and accelerating discoveries.
@@ 2 Related Work
Artificial intelligence (AI) has improved biomedical applications for processing different textual data [31, 32, 33]. Conventionally, small-scale language models [34] were used, but scaling laws indicate that increasing model parameters brings enhanced performance, leading to superior reasoning capability and higher answer accuracy. Therefore, we have seen large amounts of applications based on large language models (LLMs) [35, 36, 37, 38, 39].
@@
Biomedical LLMs. Large language models (LLMs) are generally pre-trained on vast open-domain corpora but lack domain-specific knowledge. To enhance domain-specific performance, several techniques are used: (1) Fine-tuning optimizes the total or a part of parameters to improve the model performance on a small, specific dataset [35, 37, 38, 40]. (2) Reinforcement learning with human feedback (RLHF) or AI feedback (RLAIF) updates the model parameters to align LLM’s responses with responses from humans or a teacher-model via a reinforcement learning framework [36, 39, 41]. (3) Prompt engineering [42] involves giving instructions or examples to LLM to enforce rules or enhance reasoning [43, 44, 45]. These methods mainly focus on question-answering (QA) tasks, like answering medical consultations [36], taking medical examinations [35] and summarizing clinical reports [33]. However, LLMs still struggle with complex tasks that require more profound understanding and reasoning[22]. Compared with QA tasks, responses for scientific research tasks are much longer and more professional and involve intricate, long logical chains. Fine-tuning and RLHF are infeasible due to the lack of instruction data, and simple prompt engineering cannot empower LLM with deep logical reasoning for professional topics.
@@
LLM-based Agents for Research. LLM-based agents offer significant advantages over the direct use of LLMs, such as actively acquiring information, interacting with environments, and stronger reasoning and planning [22]. LLMs can function as a single agent by being assigned with specific roles through in-domain fine-tuning [29] or role-specific prompts [33, 46, 47, 48, 49, 50, 51]. In contrast, a Multi-Agent System (MAS) comprises multiple LLM-based agents, enabling task completion through various cooperative methods. One approach involves assigning distinct roles to agents, who then reach a consensus through negotiation, like discussing clinical diagnosis via several medical experts [52, 53, 54]. However, the consensus can be unreliable due to potential instability and hallucinations from LLMs. Alternatively, MAS can distribute complex tasks into sub-tasks among agents, such as dividing a scientific discovery process into creating ideas, experimentation, and writing [28]. MAS is promising for tasks with complex logical chains and highly detailed requirements. After decomposing tasks and assigning them to agents, we can synthesize the results effectively. Furthermore, employing a professional and rigorous workflow framework helps constrain the agents’ generative processes, ensuring highly reliable outcomes.
@@
AI for Research. AI for Research (AI4R) [55] has gained significant attention, particularly in automating scientific workflows. Existing studies are categorized into four automation levels [22]: Level 0 automation performs specific predefined tasks [56]; Level 1 automation designs simple experimental protocols with in-silico or lab tools [46, 57]; Level 2 automation develops rigorous experimental protocols and employs statistical methods for hypothesis evaluation [28, 58]. Level 3 agents, which remain undeveloped, are envisioned to discover new methods and employ diverse techniques to measure biological phenomena.
@@
Our work belongs to Level 2 automation and differs from current AI4R systems. (1) The system is not limited to solving a specific biomedical task. For example, CRISPR-GPT [59] customizes the workflow only for gene editing experiments. (2) We automate the entire process without manual maintenance of a template database like Genesis [60]. (3) Most level 2 AI4R systems are designed for computer science (CS) [28, 29, 30, 61, 62], where public datasets like OpenReview are available, facilitating training and evaluation. Such extensive review data is lacking in the biomedical domain. (4) Existing studies emphasize novelty while we focus on reliability and feasibility, necessitating fine-grained literature analysis. For example, AIScientist [28] improves existing solutions based on a given task and initial experimental code. Conversely, BioResearcher designs a series of executable experiments for a new research subject.
@@ 3 BioResearcher
Figure 2: The flowchart of BioResearcher. On the left, the system is divided into four main modules: Search, Literature Processing, Experimental Design, and Programming. The right side illustrates the system’s workflow: it initiates with the Search module that retrieves pertinent literature and datasets based on the user input. The Literature Processing module then converts the obtained literature into standardized experimental reports, analyzing them against the research objective, conditions, and requirements. It also works with the Search module to evaluate the applicability of the datasets mentioned in the reports. These processed reports, analyses, and suitable datasets are then sent to the Experimental Design module, which designs an experimental protocol aligned with the research objective. Finally, the Programming module extracts the dry lab experimental tasks from the protocol and generates executable code to perform these tasks. Components underlined are agents based on LLMs.
@@ 3.1 Framework Overview
To emulate the workflow of human researchers, BioResearcher consists of four primary modules: Search, Literature Processing, Experimental Design, and Programming. As shown in Figure 2, the research process begins with the Search module, which comprehends the user input, generates appropriate queries, and searches for relevant research papers and datasets from online repositories. The retrieved literature is then filtered, downloaded, and forwarded to the Literature Processing module. Here, each research paper is standardized into an experimental report, and each report is analyzed in light of the user-specified objective, conditions, and requirements. This module also interacts with the Search module to identify the usability of the datasets mentioned in the reports. The processed reports, analyses, and datasets are then forwarded to the Experimental Design module, which constructs an experimental protocol. Finally, the Programming module extracts a sequence of dry lab experiment tasks from the protocol and generates accurate and executable codes for these tasks.